# FaithSCAN: Model-Driven Single-Pass Hallucination Detection for Faithful Visual Question Answering

### v1

@@ I Introduction
Visual Question Answering (VQA) is a representative image-to-text (I2T) task in which a model generates natural language answers given an image and a corresponding question [huang2025survey]. Recent advances in multimodal large language models (MLLMs) and vision-language models (VLMs) have enabled impressive integration of visual perception and language understanding, supporting increasingly complex open-ended VQA tasks [ging2024openended, Yu2024TowardsOV]. Despite impressive performance, they frequently generate fluent yet ungrounded or incorrect responses, a phenomenon known as hallucination [ji2023survey, farquhar2024detecting, bai2024hallucination], which in VQA settings manifests as responses misaligned with the input visual evidence or contradictory to factual knowledge [liu2024survey, hu2024mitigating]. Among them, faithfulness hallucinations [huang2025survey], where answers violate visual grounding, are particularly critical in VQA, since reliable factual verification depends on the model first attending to and correctly interpreting the image [li-etal-2023-evaluating]. Consequently, detecting faithfulness hallucinations is a necessary step for ensuring answer reliability and user trust, motivating the development of effective detection methods that can identify unfaithful answers (exmaples illustrated in Fig. 1).
Motivated by the importance of faithfulness hallucination detection in VQA, a variety of detection methods have been proposed in recent years. Existing approaches can be broadly grouped into two paradigms: external verification-based methods and uncertainty-driven methods [bai2024hallucination, chen2025survey]. External verification approaches detect hallucinations by explicitly checking cross-modal consistency between the generated answer and the visual input, often relying on auxiliary components such as object detectors, VQA models, scene graphs, or large language model (LLM) judges [li-etal-2023-evaluating, jing2024faithscore, gu2024survey]. By decomposing answers into atomic claims and validating them against visual evidence, these methods are capable of identifying fine-grained inconsistencies. However, their reliance on additional pretrained models and multi-stage pipelines leads to high computational overhead and makes the final performance sensitive to error propagation from external components. Moreover, such methods operate largely independently of the internal reasoning process of the target VLM, limiting their scalability and integration in real-world deployment [jing2024faithscore].
@@
In contrast, uncertainty-driven methods approach hallucination detection from the perspective of model confidence. Empirical studies in both language models and vision–language models have shown that hallucinated outputs are often associated with elevated predictive uncertainty [huang2025survey, liu2024survey]. Representative approaches, such as semantic entropy (SE) [farquhar2024detecting] and likelihood-based measures [farquhar2024detecting, zhang2024vl, ding2024hallu, tong2025semantic], estimate hallucination likelihood through sampling model outputs under different decoding strategies, prompt reformulations, or visual perturbations, followed by entropy or distributional analysis. Recent work enhances uncertainty estimation by incorporating visual cues, such as vision-amplified semantic entropy [liao2025vision] and dense geometric entropy [gautam2025hedge]. In addition, methods leveraging internal signals like attention patterns and intermediate embeddings are emerging. Approaches such as DHCP [zhang2025dhcp] and OPERA [huang2024opera] focus on attention dynamics to detect over-reliance on unverified features, while HallE-Control [zhai2023halle] modifies embeddings to suppress unreliable visual details. These methods highlight the value of integrating internal uncertainty signals.
@@ II-A Hallucination Detection in Visual Question Answering
Faithfulness hallucination detection is critical in VQA, where generated answers can be fluent and confident yet inconsistent with visual evidence [chen2025survey]. A variety of detection methods have been proposed in recent years, which can be broadly categorized into external verification-based and internal signal- or uncertainty-driven approaches. External verification methods explicitly check cross-modal consistency between answers and visual inputs, often using object detectors, VQA models, scene graphs, or LLM evaluators [li-etal-2023-evaluating, jing2024faithscore, gu2024survey]. By decomposing answers into atomic claims and validating them against visual evidence, they enable fine-grained hallucination detection. However, these methods rely on multi-stage pipelines and pretrained components, incurring high computational cost and sensitivity to error propagation, while largely ignoring the VLM’s internal reasoning process [jing2024faithscore]. Internal signal- or uncertainty-driven methods, in contrast, exploit the model’s outputs, attention patterns, and intermediate embeddings to detect hallucinations, offering a more direct and efficient view into the generation process [liu2025paying, jiang2025devils, ghosh2025visual].
@@ II-B Uncertainty-Based Hallucination Detection
Model-internal uncertainty has emerged as an effective and relatively inexpensive cue for hallucination detection. Both language models and VLMs exhibit increased predictive uncertainty when hallucinations occur [zhang2023enhancing, quevedo2024detecting, aryal2025howard, gu2024survey]. Methods such as semantic entropy (SE) [farquhar2024detecting] and likelihood-based measures [ma2025semantic, zhao2025sese, jiang2025interpreting] quantify uncertainty by aggregating variations across multiple decoding outputs. In VQA, extensions incorporate visual perturbations or multimodal semantic clustering, e.g., vision-amplified semantic entropy [liao2025vision] and dense geometric entropy [gautam2025hedge], which estimate hallucination likelihood by comparing semantic distributions under clean and perturbed visual inputs. These methods demonstrate that uncertainty signals can effectively capture hallucinations, including in safety-critical domains such as medical VQA [liao2025vision]. Nevertheless, most approaches rely on repeated sampling or perturbations to obtain reliable estimates, which increases inference cost and can be sensitive to stochastic decoding behavior.
@@ II-C Sources and Exploitation of Internal Uncertainty Signals
Hallucinations in VQA are not only reflected in the magnitude of predictive uncertainty but also in structured patterns across attention, intermediate representations, and logits. Early studies leveraged attention dynamics to detect hallucinations, e.g., DHCP [zhang2025dhcp] captures cross-modal attention patterns, showing consistent differences between hallucinated and faithful responses. OPERA [huang2024opera] demonstrates that hallucinations correlate with over-reliance on sparse summary tokens in self-attention, which can be mitigated through dynamic reweighting. Feature-based methods, such as HallE-Control [zhai2023halle], reduce hallucinations by projecting intermediate embeddings to suppress reliance on unverifiable visual details. Embedding- or logit-based approaches, e.g., PROJECTAWAY [jiang2025interpreting] and ContextualLens [phukan2025beyond], exploit token-level predictions and context-aware embeddings to localize and mitigate hallucinations at fine-grained object or patch levels. Despite their effectiveness, these methods are often limited by partial access to internal signals or simplified modeling of cross-modal interactions. Together, they indicate that hallucination-related cues are encoded across heterogeneous internal signals, highlighting the feasibility and necessity of fully exploring and exploiting them for reliable detection.
More recently, HaluNet [tong2025halunet] studies efficient hallucination detection in text-based question answering using internal uncertainty signals. Our work instead focuses on faithfulness hallucinations in VQA and extends this paradigm through explicit fusion of multimodal internal evidence. In addition, we provide a more systematic design and empirical validation of this approach in the multimodal setting.
@@ III-B Faithfulness Hallucination Detection in VQA
We focus on faithfulness hallucinations [huang2025survey], where the generated answer is inconsistent with the visual evidence present in the input image. Given a VQA instance:
@@ III-C Vision–Language Model Architectures
III-C1 InstructBLIP
InstructBLIP [dai2023instructblip] employs a query-based cross-modal alignment strategy. A frozen vision encoder first extracts dense visual features from the image. These features are then queried by a lightweight Query Transformer (Q-Former), producing a compact set of visual query embeddings. The embeddings are projected into the input space of a Flan-T5 language model, enabling instruction-following VQA without fully fine-tuning the vision backbone or language model.
III-C2 LLaVA-NeXT
LLaVA-NeXT adopts the same projection-based multimodal alignment paradigm as the original LLaVA [liu2023visual]. Its vision encoder (vision tower) extracts high-dimensional visual features, which are projected by a learnable multimodal layer and concatenated with text token embeddings. The combined embeddings are fed into a decoder-only LLaMA model, enabling cross-modal reasoning while treating visual features as additional tokens aligned to the language space. We use LLaVA-NeXT in this work as a stronger instantiation of the LLaVA architecture.
III-C3 Qwen-VL
Qwen3-VL[Qwen3-VL] adopts a unified multimodal transformer architecture. Visual features are embedded and merged into a sequence of multimodal tokens, which are processed alongside textual tokens across multiple transformer layers. Unlike projection-based models, visual and textual representations interact throughout the network, allowing richer semantic alignment. This tight integration improves cross-modal reasoning but also complicates uncertainty propagation, which is critical for hallucination detection.
III-D Implications for Hallucination Detection
Despite architectural differences, all three VLMs encode rich uncertainty information in their internal representations, including token-level predictive uncertainty, visual representation reliability, and cross-modal semantic alignment. These shared properties motivate a general hallucination detection approach that leverages internal representations and uncertainty signals, rather than relying on generation-level sampling or external verification.
@@ IV-B Why Supervised Hallucination Detection
Most uncertainty-based hallucination detection methods [farquhar2024detecting, zhang2024vl, ma2025semantic, tong2025semantic, zhao2025sese] aim to estimate or calibrate uncertainty, yet converting these scores into decisions still requires carefully chosen thresholds, limiting practical utility. By framing detection as a supervised decision-making problem, we directly map internal uncertainty signals to hallucination labels, producing actionable outputs and simplifying deployment. As described in Section IV-E, this approach learns from internal signals rather than calibrating model uncertainty, reducing learning difficulty: learning focuses on extracting useful internal signals rather than output-space confidence modeling.
@@ IV-C Low-Cost Supervision via Model-Driven Labeling
A central challenge in supervised hallucination detection is the acquisition of reliable supervision at scale. While manual annotation of multimodal hallucinations is prohibitively expensive, many existing benchmarks provide only answer-level correctness or predefined hallucination references [goyal2017vqa, li-etal-2023-evaluating, wang2024haloquest, park2025halloc]. Similar research has tackled this challenge [chen2024mllm], however, hallucination annotations generated by these methods are typically model-agnostic, focusing on generic failure cases without considering that hallucinations are inherently model-dependent phenomena. Since hallucinations arise from specific generation behaviors, inductive biases, and cross-modal alignment failures of the underlying vision–language model, decoupling hallucination labels from the base model limits their practical utility for detection.
In contrast, we argue that reliable supervision for hallucination detection should be model-aware, grounded in the actual failure modes exhibited during inference. To this end, we propose a low-cost, model-driven supervision strategy inspired by the LLM-as-a-Judge paradigm [gu2024survey, chen2024mllm], and extend it to the multimodal hallucination setting, as shown in Fig. 2b. The core idea is to leverage a strong vision–language model as a semantic judge to evaluate whether a candidate answer is faithfully grounded in the visual input and consistent with the question–answer pair, with the image serving as explicit evidence.
Unlike previous work, our approach incorporates a novel model-driven framework by explicitly integrating visual grounding constraints through a structured Visual-NLI (Visual Natural Language Inference) protocol. This allows hallucination labels to be tightly coupled with the model’s internal state and generation behavior. Importantly, our goal is not to introduce an additional black-box oracle for hallucination detection, but rather to design the supervision process to remain fundamentally tied to the underlying model’s inference mechanics. Specifically, the supervision signal is derived from two complementary aspects: (i) hallucination labels are conditioned on semantic consistency judgments that align with the same image–question context used by the target model, directly linking the labels to the model’s internal reasoning processes; and (ii) the supervision process is coupled with the model’s generation behavior, ensuring that the generated hallucination labels reflect the specific failure modes exhibited by the model during inference.
To make this supervision signal reliable, we employ two targeted strategies. The following sections describe these components in detail.
IV-C1 Answer-Level Model-Driven Hallucination Labels with Gold References
@@ IV-C2 Human Verification via Small-Scale Multi-Round Sampling
Concretely, we randomly sample a subset of VQA instances (100 samples per data set) and inspect the most-likely answer generated by the target VLM under low-temperature decoding. For each sampled instance, human annotators are presented with the image, question, gold answer, and model-generated answer, and must determine whether the model answer is a hallucination—defined as factually incorrect or unsupported by the visual evidence. To avoid confounding factors, samples with ambiguous questions or problematic ground-truth references are explicitly marked and excluded from subsequent agreement analysis.
Importantly, human annotation is not used to replace model-driven labels. Instead, it is employed to validate whether the hallucination signals produced by the Visual-NLI judge are consistent with human judgment when applied to the same model outputs. Let $y^{\text{human}}\in\{0,1\}$ denote the human label indicating non-hallucination or hallucination, and $y^{\text{model}}\in\{0,1\}$ denote the corresponding binary label derived from the model-driven NLI judgment. We evaluate the agreement between $y^{\text{human}}$ and $y^{\text{model}}$ using multiple complementary metrics, including raw agreement ratio, Cohen’s $\kappa$, and Matthews correlation coefficient (MCC); detailed analyses and discussions are provided in Section VI-B.
This verification process serves two purposes. First, it empirically assesses the reliability of the model-driven supervision signal without requiring large-scale manual annotation. Second, it provides evidence that hallucination labels constructed via Visual-NLI judgments are not arbitrary artifacts of the judge model, but are aligned with human perception of visual faithfulness. As such, the human verification step supports the use of model-driven labels as effective supervision for training hallucination detectors.
@@ IV-D2 Token-Level Generation Signals
At each decoding step $t$, the language model outputs a predictive distribution
@@ IV-D3 Visual Semantic Representations
Prior to cross-modal interaction, the vision encoder extracts visual features from the input image $I$. We denote the resulting set of patch-level visual embeddings as
@@ IV-D5 Unified View of Uncertainty Signals
IV-D6 Unified View of Uncertainty Signals
Although the concrete extraction of the above signals is tightly coupled with the architecture of a given VLM, their roles are conceptually unified. Token-level generation signals capture uncertainty in language decoding, visual semantic representations reflect uncertainty in visual understanding, and aligned multimodal representations encode uncertainty arising from vision–language interaction.
@@ IV-E2 Branch-wise Evidence Encoding
All branch encoders are implemented as lightweight sequence compression modules rather than deep semantic encoders. Specifically, we consider linear projections with global average pooling, trainable sequence compressors (Linear + LayerNorm + ReLU + mean pooling), and convolutional encoders with adaptive pooling [kim2014convolutional], where one-dimensional convolutions are applied with kernel size $3$ and stride $1$. Despite architectural differences, all variants share the same objective: to summarize the reliability and stability of internal signals while limiting representational capacity to reduce overfitting. We further analyze the generalization capability under distribution shifts in Section VI-A2, while the effectiveness of architectural variants is examined in Section VI-C2.

### v2

@@ I Introduction
Visual Question Answering (VQA) is a representative image-to-text (I2T) task in which a model generates natural language answers given an image and a corresponding question [19]. Recent advances in multimodal large language models (MLLMs) and vision-language models (VLMs) have enabled impressive integration of visual perception and language understanding, supporting increasingly complex open-ended VQA tasks [12, 43]. Despite impressive performance, they frequently generate fluent yet ungrounded or incorrect responses, a phenomenon known as hallucination [21, 9, 3], which in VQA settings manifests as responses misaligned with the input visual evidence or contradictory to factual knowledge [29, 18]. Among them, faithfulness hallucinations [19], where answers violate visual grounding, are particularly critical in VQA, since reliable factual verification depends on the model first attending to and correctly interpreting the image [27]. Consequently, detecting faithfulness hallucinations is a necessary step for ensuring answer reliability and user trust, motivating the development of effective detection methods that can identify unfaithful answers (exmaples illustrated in Fig. 1).
Motivated by the importance of faithfulness hallucination detection in VQA, a variety of detection methods have been proposed in recent years. Existing approaches can be broadly grouped into two paradigms: external verification-based methods and uncertainty-driven methods [3, 5]. External verification approaches detect hallucinations by explicitly checking cross-modal consistency between the generated answer and the visual input, often relying on auxiliary components such as object detectors, VQA models, scene graphs, or large language model (LLM) judges [27, 24, 15]. By decomposing answers into atomic claims and validating them against visual evidence, these methods are capable of identifying fine-grained inconsistencies. However, their reliance on additional pretrained models and multi-stage pipelines leads to high computational overhead and makes the final performance sensitive to error propagation from external components. Moreover, such methods operate largely independently of the internal reasoning process of the target VLM, limiting their scalability and integration in real-world deployment [24].
@@
In contrast, uncertainty-driven methods approach hallucination detection from the perspective of model confidence. Empirical studies in both language models and vision–language models have shown that hallucinated outputs are often associated with elevated predictive uncertainty [19, 29]. Representative approaches, such as semantic entropy (SE) [9] and likelihood-based measures [9, 45, 8, 41], estimate hallucination likelihood through sampling model outputs under different decoding strategies, prompt reformulations, or visual perturbations, followed by entropy or distributional analysis. Recent work enhances uncertainty estimation by incorporating visual cues, such as vision-amplified semantic entropy [28] and dense geometric entropy [10]. In addition, methods leveraging internal signals like attention patterns and intermediate embeddings are emerging. Approaches such as DHCP [47] and OPERA [20] focus on attention dynamics to detect over-reliance on unverified features, while HallE-Control [44] modifies embeddings to suppress unreliable visual details. These methods highlight the value of integrating internal uncertainty signals.
@@ II-A Hallucination Detection in Visual Question Answering
Faithfulness hallucination detection is critical in VQA, where generated answers can be fluent and confident yet inconsistent with visual evidence [5]. A variety of detection methods have been proposed in recent years, which can be broadly categorized into external verification-based and internal signal- or uncertainty-driven approaches. External verification methods explicitly check cross-modal consistency between answers and visual inputs, often using object detectors, VQA models, scene graphs, or LLM evaluators [27, 24, 15]. By decomposing answers into atomic claims and validating them against visual evidence, they enable fine-grained hallucination detection. However, these methods rely on multi-stage pipelines and pretrained components, incurring high computational cost and sensitivity to error propagation, while largely ignoring the VLM’s internal reasoning process [24]. Internal signal- or uncertainty-driven methods, in contrast, exploit the model’s outputs, attention patterns, and intermediate embeddings to detect hallucinations, offering a more direct and efficient view into the generation process [31, 23, 11].
@@ II-B Uncertainty-Based Hallucination Detection
Model-internal uncertainty has emerged as an effective and relatively inexpensive cue for hallucination detection. Both language models and VLMs exhibit increased predictive uncertainty when hallucinations occur [46, 37, 1, 15]. Methods such as semantic entropy (SE) [9] and likelihood-based measures [32, 48, 22] quantify uncertainty by aggregating variations across multiple decoding outputs. In VQA, extensions incorporate visual perturbations or multimodal semantic clustering, e.g., vision-amplified semantic entropy [28] and dense geometric entropy [10], which estimate hallucination likelihood by comparing semantic distributions under clean and perturbed visual inputs. These methods demonstrate that uncertainty signals can effectively capture hallucinations, including in safety-critical domains such as medical VQA [28]. Nevertheless, most approaches rely on repeated sampling or perturbations to obtain reliable estimates, which increases inference cost and can be sensitive to stochastic decoding behavior.
@@ II-C Sources and Exploitation of Internal Uncertainty Signals
Hallucinations in VQA are not only reflected in the magnitude of predictive uncertainty but also in structured patterns across attention, intermediate representations, and logits. Early studies leveraged attention dynamics to detect hallucinations, e.g., DHCP [47] captures cross-modal attention patterns, showing consistent differences between hallucinated and faithful responses. OPERA [20] demonstrates that hallucinations correlate with over-reliance on sparse summary tokens in self-attention, which can be mitigated through dynamic reweighting. Feature-based methods, such as HallE-Control [44], reduce hallucinations by projecting intermediate embeddings to suppress reliance on unverifiable visual details. Embedding- or logit-based approaches, e.g., PROJECTAWAY [22] and ContextualLens [36], exploit token-level predictions and context-aware embeddings to localize and mitigate hallucinations at fine-grained object or patch levels. Despite their effectiveness, these methods are often limited by partial access to internal signals or simplified modeling of cross-modal interactions. Together, they indicate that hallucination-related cues are encoded across heterogeneous internal signals, highlighting the feasibility and necessity of fully exploring and exploiting them for reliable detection.
More recently, HaluNet [40] studies efficient hallucination detection in text-based question answering using internal uncertainty signals. Our work instead focuses on faithfulness hallucinations in VQA and extends this paradigm through explicit fusion of multimodal internal evidence. In addition, we provide a more systematic design and empirical validation of this methodology in the multimodal setting.
@@ III-B Faithfulness Hallucination Detection in VQA
We focus on faithfulness hallucinations [19], where the generated answer is inconsistent with the visual evidence present in the input image. Given a VQA instance:
@@ III-C Vision–Language Model Architectures
Existing vision–language models vary in how they align visual and textual modalities. We focus on three representative models that illustrate distinct paradigms.
  • InstructBLIP [7] uses query-based alignment, producing compact visual embeddings that interact with the language model via queries.
  • LLaVA-NeXT [30] relies on projection-based alignment, treating visual features as additional text tokens concatenated with language embeddings.
  • Qwen-VL[2] employs a unified multimodal transformer, where visual and textual tokens interact throughout all layers.
Despite these architectural differences, all models encode rich uncertainty information, including token-level predictive uncertainty, visual reliability, and cross-modal alignment. This motivates hallucination detection methods that leverage internal representations rather than relying solely on generation-level sampling. Details of each architecture are provided in Supplementary Material Sec. B.1.
@@ IV-B Why Supervised Hallucination Detection
Most uncertainty-based hallucination detection methods [9, 45, 32, 41, 48] aim to estimate or calibrate uncertainty, yet converting these scores into decisions still requires carefully chosen thresholds, limiting practical utility. By framing detection as a supervised decision-making problem, we directly map internal uncertainty signals to hallucination labels, producing actionable outputs and simplifying deployment. As described in Section IV-E, this approach learns from internal signals rather than calibrating model uncertainty, reducing learning difficulty: learning focuses on extracting useful internal signals rather than output-space confidence modeling.
@@ IV-C Low-Cost Supervision via Model-Driven Labeling
A central challenge in supervised hallucination detection is obtaining reliable supervision at scale. Manual annotation of multimodal hallucinations is costly, and many benchmarks provide only correct answers or predefined hallucination references [13, 27, 42, 34]. Automatic annotation methods can alleviate this [4], but they typically produce model-agnostic labels that capture generic failures rather than the model-dependent nature of hallucinations, which stem from a model’s generation behaviors, inductive biases, and cross-modal alignment deficiencies. Consequently, supervision decoupled from the target model is inherently limited.
We generate hallucination labels tightly coupled with the target model’s decoding dynamics [13, 27, 42, 34], derived from semantic consistency between generated and reference answers conditioned on the same image–question context. A strong external VLM serves as a semantic judge [15, 4], but it is not a detached oracle; it evaluates answers in the context of the target model’s reasoning, producing model-aware labels. Visual grounding constraints enforced via a structured Visual-NLI protocol ensure that labels capture model-specific failure modes. We then verify label reliability through complementary automatic and human-based procedures, which is critical for supervised hallucination detection.
IV-C1 Answer-Level Hallucination Labels Conditioned on Gold Answers
@@ IV-C2 Human Verification via Small-Scale Multi-Round Sampling
Concretely, we randomly sample a subset of VQA instances (100 samples per data set) and inspect the most-likely answer generated by the target VLM under low-temperature decoding. For each sampled instance, human annotators are presented with the image, question, gold answer, and model-generated answer, and must determine whether the model answer constitutes a hallucination, defined as being unfaithful to the visual evidence or inconsistent with the reference answer. To avoid confounding factors, samples with ambiguous questions or problematic ground-truth references are explicitly marked and excluded from subsequent agreement analysis.
Importantly, human annotation is not used to replace model-driven labels. Instead, it is employed to validate whether the hallucination signals produced by the Visual-NLI judge are consistent with human judgment when applied to the same model outputs. Let $y^{\text{human}}\in\{0,1\}$ denote the human label indicating non-hallucination or hallucination, and $y^{\text{model}}\in\{0,1\}$ denote the corresponding binary label derived from the model-driven NLI judgment. We evaluate the agreement between $y^{\text{human}}$ and $y^{\text{model}}$ using multiple complementary metrics, including raw agreement ratio, Cohen’s $\kappa$, and Matthews correlation coefficient (MCC); detailed analyses and discussion are provided in Section VI-B.
IV-C3 Automated Label Reliability Verification
Assessing the reliability of such labels through human verification remains costly and inherently limited. We therefore introduce an automated label reliability verification framework that aims to quantify the reliability of model-driven hallucination labels in a scalable manner, and to examine how such reliability estimates can be incorporated into the training process.
Given an image–question–answer triplet, the verification procedure evaluates the consistency of the assigned hallucination label across multiple model-based assessments, including repeated semantic verification [9], stochastic answer generation [41], and explicit self-reflection. These assessments capture complementary aspects of decisiveness, robustness, and self-consistency, and are aggregated into an instance-level reliability score that reflects confidence in the supervision signal rather than the likelihood of hallucination itself.
We further investigate the use of this reliability score as an uncertainty-aware signal to modulate the contribution of individual training instances, without altering the original labels or introducing inference-time overhead. Detailed definitions of the methods, strategies for training integration, and preliminary experimental validations are provided in Supplementary Material Sections D.2, D.3.
@@ IV-D2 Token-Level Generation Signals
At each decoding step $t$, the language model outputs a predictive distribution:
@@ IV-D3 Visual Semantic Representations
Prior to cross-modal interaction, the vision encoder extracts visual features from the input image $I$. We denote the resulting set of patch-level visual embeddings as:
@@ IV-E2 Branch-wise Evidence Encoding
All branch encoders are implemented as lightweight sequence compression modules rather than deep semantic encoders. Specifically, we consider linear projections with global average pooling, trainable sequence compressors (Linear + LayerNorm + ReLU + mean pooling), and convolutional encoders with adaptive pooling [25], where one-dimensional convolutions are applied with kernel size $3$ and stride $1$. For a detailed description of these branch encoders, refer to Supplementary Material Sec. B.5. Despite architectural differences, all variants share the same objective: to summarize the reliability and stability of internal signals while limiting representational capacity to reduce overfitting. We further analyze the generalization capability under distribution shifts in Section VI-A2, while the effectiveness of architectural variants is examined in Section VI-C2.